Integrate the following in terms of dx:

(x^2 – 2x +3) / x^4

Grade:11

2 Answers

Vikas TU
7 years ago
split (x^2 – 2x +3) / x^4 into:

(x^-2 -2x^-3 + 3x^-4) dx
=  -x^-1 + x^-2 – x^-3 + c
= -1/x + 1/x^2 – 1/x^3 + c

 
SREEKANTH
7 years ago
integtration of =(x^2 – 2x +3) / x^4
                   =(x^-2-2x^-3+3x^-4)
                   =integration of x^-2 dx-integration of -2x^-3 dx+integration of 3x^-4
                   =(x^-2+1/-2+1)-(2x^-3+1/-3+1)+(3x^-4+1/-4+1)
                   =-1/x+1/x^2-1/x^3+c
         where c is an arbitrary constant